OpenStudy (anonymous):

what is the area of the region bounded below by y=e^x and above by y=4e^-x, for x is greater than or equal to zero?

OpenStudy (anonymous):

first find the point at which the upper and lower boundaries intersect. e^x=4 e^(-x) x=ln(4)-x 2x=ln(4) x=ln(2) next integrate to find the area integrate(4e^(-x)-e^x,{x,0,ln(2)})

OpenStudy (anonymous):

thank you :)